OpenStudy (anonymous):

I need help simplifying the expression 11^2/5 / 11^4/5 into (the fifth root of 1331) / 11 ...the 5th square root applies only to the 1331 in the numerator! the 11 as the denominator is just an 11! please help!

OpenStudy (pfenn1):

Is this your equation?\[{\frac{11^2}{5 }\over \frac{11^4}{5}}\div \frac{\sqrt[5]{1331}}{11}\]

OpenStudy (anonymous):

where you have the division sign...it should be an equals sign

OpenStudy (pfenn1):

Like this?\[{\frac{11^2}{5 }\over \frac{11^4}{5}}= \frac{\sqrt[5]{1331}}{11}\]

OpenStudy (anonymous):

yes, that is the correct setup

OpenStudy (pfenn1):

\[{\frac{11^2}{5 }\ \times \frac{5}{11^4}{}}= \frac{\sqrt[5]{1331}}{11}\]Can you simplify it from here?

OpenStudy (anonymous):

Im not sure how what is on the left side of the "=" sign can equal what is on the right!

OpenStudy (pfenn1):

I was wondering that as well. Are you sure that you copied your original equation correctly?\[{\frac{11^2}{5 }\over \frac{11^4}{5}}= \frac{\sqrt[5]{1331}}{11}\]

OpenStudy (anonymous):

yea...the answer, on the right of the "=" sign, is what the book says is the answer

OpenStudy (pfenn1):

The answer to what?

OpenStudy (anonymous):

the left side (the problem) is equal to the right side (the answer) ....Im not sure why the answer is the answer!